page 389
for each pair funtion,find fog and gof , if they are exit
24)
f = {(3,8),(4,0),(6,3),(7,-1)}
g = {(0,4),(8,6)(3,6),(-1,8)}
26)
f = {(2,5),(3,9),(-4,1)}
g = {(5,-4),(8,3),(2,-2)}

Answer by stanbon(75887)

for each pair funtion,find fog and gof , if they are exist
24)
f = {(3,8),(4,0),(6,3),(7,-1)}
g = {(0,4),(8,6)(3,6),(-1,8)}
---
fog(x) = f[g(x)]
---
fog(0) = f(4) =0
fog(8) = f(6) = 3
fog(3) = f(6) = 3
fog(-1) =f(8) = not defined
---
fog(x) is a function on {0,8,3}
--------------------------------
gof(3) = g(8) =6
gof(4) = g(0) = 4
gof(6) = g(3) = 6
gof(7) = g(-1) = 8
gof(x) is a function on {3,4,6,7}
=======================================
Cheers,
Stan H.
====================
=================================


26)
f = {(2,5),(3,9),(-4,1)}
g = {(5,-4),(8,3),(2,-2)}